OpenStudy (anonymous):

A semicircular window of radius 12 inches is to be laminated with a sunblock coating that costs $0.80 per square inch to apply. What is the total cost of coating the window, to the nearest cent? (Use p = 3.14. Hint: Area of Circle: A = p r²

OpenStudy (ybarrap):

Area = pi r^2 = 3.14 * 12^2 = 452.16 sq/in Total Cost = Area * $.80/sq/in = 452.16 sq/in * $.80/sq/in = $361.73 Ans. $361.73

OpenStudy (anonymous):

I have been struggling with this problem all day and those answers never came up, wow. thnx man